A contractor purchased an equipment with a guarantee for 3 years so he deposited $10,000 in the bank to cover the total maintenance cost. He will start withdrawing the maintenance cost of $ 1000 by the end of the 3rd year for 5 years. What will be the remaining amount in the bank at the end of the 10th year from now assuming the bank pays 6 percent compounded annually?
Answer #1

rate positively ..

i ii iii=i*6% iv=i-ii+iii
Year Beginning value Withdraw Interest earned ending balance
0        10,000.00
1          10,000.00             600.00        10,600.00
2          10,600.00             636.00        11,236.00
3          11,236.00    1,000.00             674.16        10,910.16
4          10,910.16    1,000.00             654.61        10,564.77
5          10,564.77    1,000.00             633.89        10,198.66
6          10,198.66    1,000.00             611.92          9,810.58
7            9,810.58    1,000.00             588.63          9,399.21
8            9,399.21             563.95          9,963.16
9            9,963.16             597.79        10,560.95
10          10,560.95             633.66        11,194.61
Answer =        11,194.61